Assignment-07: Problem 1 Previous Problem Problem List Next Problem (8 points) The wait time (after a scheduled arrival time)Please show all work. Thank you!
Answer #1

Here x ~ Uco, 15) , n=95. we want p(670 L EX <796), Now we know that X= {X/n. 1. divide it by n we get, P (670 <EXi <796) =PE = √18.75 = 4.33 p(7.05 < x < 8.33) = P17.05-7.5 T 4.33 95 L X-1 28.33-7.01 75 4331895 = P(-1.01 <z < 1.87) = P(Z <1.87) - -Data P147,92) = I-PCV 291) 1 - 885588 = 0.114412 obtained using R software pbinom (91,950.9332) d p (x>5) - PCX- 5-7.5 4.33)